Compute the probability of x successes using the binomial formula. Round your answers to three decimal places as needed.

n = 3, p = 0.04, X = 1
Answer #1

Solution :

Given that ,

p = 0.04

1 - p = 1 -0.04 = 0.96  

n = 3

Using binomial probability formula ,

P(X = x) = ((n! / (n - x)!) * px * (1 - p)n - x

P(X = 1) = ((3! / (3 - 1)!) * 0.041 * 0.963 - 1

=0.1106

Probability =0.1106